MULTIPLE CHOICE

A company has outstanding debt with a market value of​ $250M and common stock with a market value of​ $550M. If its debt has a​ before-tax cost of​ 7%, a​ before-tax cost of equity of​ 10% and a corporate tax rate of​ 40%, what is its​ WACC?

A.​7.86%
B.​5.44%
C.​6.55%
D.8.19%
Answer #1

After-tax cost of debt=7*(1-tax rate)

=7*(1-0.4)=4.2%

Total value=(250+550)=$800 M

WACC=Respective costs*Respective weight

=(250/800*4.2)+(550/800*10)

which is equal to

=8.19%(Approx).